OpenStudy (anonymous):

Chad is tied to one end of a 40-m elasticized (bungee) cord. The other end of the cord is tied to the middle of a bridge. If Chad jumps off the bridge, for how long will he fall before the cord begins to stretch? (Use 4.9t2 = s.)

OpenStudy (anonymous):

equation of motion: y = 0.5 a t^2 + Vot + yo y = 0.5 a t^2 + Vot + yo with yo=0, Vo =0, y = -40 and a = -10 -40 = 0.5 (-10) (t^2) + 0t + 0 -40 = 0.5 (-10) (t^2) 4 = 0.5 (t^2) 8 = t^2 t = sqrt(8) SIMPLIFIED ANSWER: 2sqrt(2)